OpenStudy (anonymous):

Solve the equation symbolically using the elimination method 12x+2y=12 =12x+3y=-12

OpenStudy (anonymous):

You can use: 12x + 2y = 12 *(-1) 12x + 3y= -12 -> -12x-2y=-12 12x+3y=-12 -> 0x+ y = -24 y=-24 and x = ( 12x - 48 = 12 => x=60/12 => x = 5)